Colgate-Palmolive Company has just paid an annual dividend of $ 1.88. Analysts are predicting dividends to grow by $ 0.12 per year over the next five years. After​ then, Colgate's earnings are expected to grow 5.9 % per​ year, and its dividend payout rate will remain constant. If​ Colgate's equity cost of capital is 7.6 % per​ year, what price does the​ dividend-discount model predict Colgate stock should sell for​ today?

The price per share is ​$ ( ) (Round to two decimal​ places.)
Answer #1
Year Dividend
1 (1.88+0.12)=2
2 (2+0.12)=2.12
3 (2.12+0.12)=2.24
4 (2.24+0.12)=2.36
5 (2.36+0.12)=2.48

Value after year 5=(D5*Growth rate)/(cost of capital-Growth rate)

=(2.48*1.059)/(0.076-0.059)

=$154.4894118

Hence current price=Future dividends*Present value of discounting factor(rate%,time period)

=2/1.076+2.12/1.076^2+2.24/1.076^3+2.36/1.076^4+2.48/1.076^5+154.4894118/1.076^5

which is equal to

=$116.08(Approx).